OpenStudy (anonymous):

How do I simplify this expression: 2√5 / 5√2? I got 2√10 / 10 but the answer is √10 / 5... how?

OpenStudy (anonymous):

I tried 2√5 / 5√2 = 2√5 / 5√2 = √2 / √2

OpenStudy (anonymous):

don't worry, 2sqrt(10) / 10 = sqrt(10)/5

OpenStudy (anonymous):

you just simplify

OpenStudy (anonymous):

what?

OpenStudy (king):

see..in the top 2 can be written as sqrt2*sqrt2 and in the bottom 5 is sqrt 5*sqrt5 so u wil get sqrt2/sqrt5 now multiply by sqrt5/sqrt5 to get sqrt10/5

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Don't I do this: 2√5 2√5 √2 ---- = ----- x ---- 5√2 5√2 √2

OpenStudy (anonymous):

yes, called rationalizing the denominator

OpenStudy (anonymous):

yes. But I got 2√10 / 10 D:

OpenStudy (anonymous):

when it's supposed to be √10 / 5

OpenStudy (anonymous):

\[\frac{2\sqrt{10}}{5\sqrt{2}*\sqrt{2}}\]

OpenStudy (anonymous):

can you see from that?

OpenStudy (anonymous):

\[= \frac{2\sqrt{10}}{5*2}\]

OpenStudy (anonymous):

\[= \frac{\sqrt{10}}{5}\]